OpenStudy (kaylaprincess):

Please help! I have an exam at nine and I don't understand how to solve this..

OpenStudy (kaylaprincess):

This is off my study guide. The question asks which of the expressions are correct. I see the answer but I don't know how they got it... I need step by step process to see!!

OpenStudy (kaylaprincess):

@Nnesha @Jaynator495 @Abhisar @ganeshie8 i have an exam soom, can you show me these steps to solve this??

Nnesha (nnesha):

f(x) is same as y so we can replace f(x) with y \[\huge\rm \color{reD}{y}=\frac{ 3\color{blue}{x}-4 }{ 5 }\] now to find inverse(f^{-1}(x) ) you should `switch x and y ` \[\huge\rm \color{blue}{x}=\frac{ 3\color{red}{y}-4 }{ 5 }\] now solve for y

OpenStudy (kaylaprincess):

multiply by five to get rid of the fraction?

Nnesha (nnesha):

yes right

OpenStudy (kaylaprincess):

doing the work one moment

OpenStudy (kaylaprincess):

oh so x = 3y - 4 ?

Nnesha (nnesha):

you should multiply both sides one u do something to the one side you Have to do the same thing to the opposite side

OpenStudy (kaylaprincess):

5x = 3y - 4 ?

Nnesha (nnesha):

yes right

OpenStudy (kaylaprincess):

add four now ?

OpenStudy (kaylaprincess):

5x + 4 = 3y ?

Nnesha (nnesha):

good right

Nnesha (nnesha):

looks good last step!

OpenStudy (kaylaprincess):

divide by three ?

Nnesha (nnesha):

that's it!

OpenStudy (kaylaprincess):

Thank you so much for your help and time! I really appreciate it.

Nnesha (nnesha):

np :=) good work!

OpenStudy (kaylaprincess):

thank you <3